# Sergei Gukov, Ludmil Katzarkov, Josef Svoboda

### v2

@@
Abstract.
We study quantum $q$-series invariants of 3-manifolds $\widehat{Z}_{\sigma}$ of Gukov–Pei–Putrov–Vafa, using techniques from the theory of normal surface singularities such as splice diagrams. We show that the (suitably normalized) sum of all $\widehat{Z}_{\sigma}$ depends only on the splice diagram, and in particular, it agrees for manifolds with the same universal abelian cover. We use these ideas to find simple formulas for $\widehat{Z}_{\sigma}$ invariants of Seifert manifolds. Applications include a better understanding of the vanishing of the $q$-series $\widehat{Z}_{\sigma}$.
Key words and phrases:
quantum invariant, 3-manifold topology, surface singularity
2020 Mathematics Subject Classification:
Primary; Secondary
1. Introduction
@@
More specifically, we study $q$-series invariants $\widehat{Z}_{\sigma}(Y;q)$ of a 3-manifold $Y$, the so-called GPPV invariants, associated with quantum groups $\mathcal{U}_{q}(\mathfrak{g})$ at generic $|q|<1$ and labeled by a $spin^{c}$ structure $\sigma\in spin^{c}(Y)$ on $Y$. They are expected to have a categorification [35].
Mostly interested in topological and geometric aspects, throughout the paper, we consider the simplest non-trivial choice of $\mathfrak{g}=\mathfrak{sl}_{2}$, which corresponds to the “gauge group” $G=SU(2)$, though much of the discussion can be generalized to higher rank root systems. [^1] We note that by “gauge group” one can mean at least two different things. One is the gauge group in complex Chern–Simons theory, for which $\widehat{Z}_{\sigma}(Y;q)$ provides a non-perturbative definition that behaves well under cutting and gluing. In contrast, the other one is a gauge group in 3d-3d correspondence and string theory fivebrane setup. The former group is a complexification of the latter, and it is the latter that we use to label $\widehat{Z}_{\sigma}(Y;q)$ in this paper. So, even when we talk about $G=SU(2)$, the gauge group in complex Chern–Simons theory and its relations to quantum groups is, in fact, $SL(2,\mathbb{C})$. Also, unless stated otherwise, throughout the paper, we work in the category of connected irreducible oriented negative definite plumbed 3-manifolds [50] with $b_{1}=0$, for which one can use the definition of $\widehat{Z}_{\sigma}(Y;q)$ as in [34] (A reader interested in various generalizations and extensions is welcome to consult [30, 2, 57, 58, 59, 7, 61].)
Our main goal is to approach the invariants $\widehat{Z}_{\sigma}(Y;q)$ by using various methods of the singularity theory of normal complex surfaces, most notably the universal abelian covers and splice diagrams. Among other things, this new perspective leads to simpler expressions for $\widehat{Z}_{\sigma}(Y;q)$ for certain families of 3-manifolds (Theorem 4.2) and sheds new light on what topological information they capture. Specifically, $\widehat{Z}_{\sigma}(Y;q)$ have close cousins $Z_{a}(Y;q)$ (without a hat!) that are not expected to admit a categorification, but are nevertheless more natural from the viewpoint of complex Chern–Simons theory [31]. While both sets of invariants can be defined independently, they are linearly related and, therefore, can be expected to contain roughly the same topological information about $Y$. By taking a closer look at the structure of the linear relation between $\widehat{Z}_{\sigma}(Y;q)$ and $Z_{a}(Y;q)$ through the looking glass of the singularity theory, we observe that $Z_{0}(Y;q)$ is, surprisingly, a much simpler invariant of $Y$ in the precise sense that we explain below.
@@
1.1. Main results
Let $Y$ be a closed, oriented, irreducible 3-manifold which is plumbed, see Section 2 for definitions. Moreover assume that the plumbing matrix can be chosen to be negative definite. We call such a 3-manifold a negative definite plumbed manifold as is traditional [^2] The reader should keep in mind that negative definiteness is a property of the plumbing graph rather than of $Y$. Note that these conditions are equivalent to $Y$ being a link of an isolated normal surface singularity $X$. Moreover, let us assume that $Y$ is a rational homology sphere, i.e., $b_{1}(Y)=0$, so it has finite $H:=H_{1}(Y,\mathbb{Z})$. Under these assumptions, $\widehat{Z}_{\sigma}(q)$ invariants were defined in [34].
Splice diagram is a combinatorial notion developed by Neumann, Eisenbud and Wahl [19, 54] for the study of singularities and 3-manifolds described above. Constructed from the plumbing graph, it contains certain essential part of it and proves very useful for studying of $\widehat{Z}_{\sigma}(q)$ invariants. The following theorem states that the series $Z_{0}(q)=\sum_{\sigma\in spin^{c}(Y)}{\widehat{Z}_{\sigma}(q)}$ essentially depends on the splice diagram (see Section 3.1 for definitions). This applies in particular to homology spheres, where it provides a useful computational tool for obtaining the unique $\widehat{Z}_{\sigma}(q)=Z_{0}(q).$ Denote by $\lambda(Y)$ the Casson–Walker invariant [70].
@@ 1.1. Main results
Topologically, splice diagrams determine universal abelian covers and vice versa [54, 60]. Universal abelian cover of a manifold $Y$ is the maximal cover of $Y$ with abelian deck group and by our assumptions, this cover is finite. We obtain the following corollary of the theorem:
@@
$q^{-6\lambda(Y_{1})}Z_{0}(Y_{1},q^{|H_{1}(Y_{1})|})=q^{-6\lambda(Y_{2})}Z_{0}(Y_{2},q^{|H_{1}(Y_{2})|}).$
For example, one can take Seifert manifolds $Y_{1}=M(1;(7,1),(7,1),(7,4))$ and $Y_{2}=M(1;(7,1),(7,2),(7,3))$ which have $Y_{1}^{ab}=Y_{2}^{ab}$ and $H=(\mathbb{Z}/7\mathbb{Z})^{2}$, and we have
$Z_{0}(Y_{1},q)=-q^{31/14}(-1+3q^{5/7}-3q^{12/7}+q^{3}-q^{11}+3q^{96/7}-\dots)=q^{18/14}Z_{0}(Y_{2},q).$
Seifert manifolds (Seifert fiber spaces) mentioned above are circle fibrations over 2-dimensional orbifolds. They admit “star-shaped” plumbing graphs and form basic building blocks of plumbed manifolds in Jaco–Shalen–Johannson decomposition. In this special case, we can go further with simplifications of $Z_{0}(Y)$ and we will give an explicit and rather simple formula, the “reduction” Theorem 4.2, using the data of Seifert fibration (rather than that of the plumbing graph). Using the group action of $H$ on $Y^{ab}$, we will extract the separate $\widehat{Z}_{\sigma}(q)$.
The next theorem gives explicit formulas for $\widehat{Z}_{\sigma}(q)$ invariants of Seifert manifolds, see Section 4 for definitions and conventions. We introduce “Laplace transform” $\mathcal{L}_{A}$, defined on monomials as
$\mathcal{L}_{A}(t^{n})=q^{n^{2}/4A}$
and extended linearly to any formal power series in $t.$ For a rational function $f(t)$, we denote $\operatorname{s.e.}f(t)$ the “symmetric expansion”, which is the average of the expansions of the given rational function as $t\rightarrow 0$ (as a Laurent power series in $t$) and as $t\rightarrow\infty$ (as a Laurent power series in $t^{-1}$). Finally, we pick a reference “anticanonical” $spin^{c}{}$ structure $\sigma^{\text{ac}}$ to express all $\widehat{Z}_{\sigma}(q)$ using the natural action of $H=H_{1}(Y,\mathbb{Z})=\langle g_{0},\dots,g_{k}\rangle$ on $spin^{c}(Y)$ (see Section 2.1 and 4 for notations and definitions).
@@
Let $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ be a Seifert manifold over $S^{2}$ ($k\geq 3$) with $e=-b+\sum_{i=1}^{k}b_{i}/a_{i}<0$. Then
(1.2) $Z_{0}(q^{|H|})=q^{\Lambda}\mathcal{L}_{A}(\operatorname{s.e.}f_{0}(t)),$
@@
(1.3) $f_{0}(t)=\frac{(t^{A_{1}}-t^{-A_{1}})(t^{A_{2}}-t^{-A_{2}})\cdots(t^{A_{k}}-t^{-A_{k}})}{(t^{A}-t^{-A})^{k-2}},$
and $\Lambda=\Lambda(Y)$ is an explicit rational number. Moreover, we have
(1.4) $\widehat{Z}_{h\sigma^{\text{ac}}}(q)=q^{\Lambda/|H|}\mathcal{L}_{A|H|}(\operatorname{s.e.}f_{h\sigma^{\text{ac}}}(t))$
@@
(1.5) $\sum_{h\in H}f_{h\sigma^{\text{ac}}}(t)h=t^{(k-2)A-\sum_{i}A_{i}}\frac{(g_{1}t^{2A_{1}}-1)\cdots(g_{k}t^{2A_{k}}-1)}{(g_{0}t^{2A}-1)^{k-2}}.$
These formulae have both computational and conceptual significance. While in the original formulation of $\widehat{Z}_{\sigma}(q)$ we need to work with the plumbing data, which can be very large due to the presence of continued fractions, here we get $\widehat{Z}_{\sigma}(q)$ simply from the data of the Seifert fibration. It also sheds light on the role of the action of $H$, leading to a better understanding of the vanishing of the $q$-series for certain $\sigma$ in Corollary 4.3. Note that the independence of $f_{0}(t)$ on $b_{i}$ in the first formula is an instance of Theorem 3.5.
1.2. Spectra in algebra and geometry
@@
The $\widehat{Z}_{\sigma}(Y)$ invariants are defined and studied using plumbing graphs. In singularity theory, plumbing graphs correspond to resolutions of the singularity which are smooth complex surfaces. Many deep results in singularity theory concern other smooth surfaces related to the singularity, namely those which are results of a smoothing. The theory of Milnor fibrations was originally developed for singularities of a single holomorphic function (hypersurface singularities) [44] and later extended to complete intersections [41].
@@
Spectrum of hypersurface singularity [68], was defined by Arnold, Steenbrink and Varchenko and generalized later by Steenbrink and Ebeling to complete intersections [18]. The spectrum is a collection of real numbers that refine the eigenvalues of the monodromy, using a natural mixed Hodge structure on the cohomology of Milnor fiber. Its most important property is the upper-semicontinuity under deformations.
@@
A relation of the Milnor fiber with the topology of the link was given by Fintushel and Stern [24]. Recall that for pairwise coprime integers $p,q,r$, Brieskorn homology sphere $Y=\Sigma(p,q,r)$ is the link of the Brieskorn singularity $X:x^{p}+y^{q}+z^{r}=0$.
Proposition 1.4 ([24]).
@@
Denote $\lambda(Y)$ be the $SU(2)$-Casson invariant of $Y$ and $\sigma(X)$ the signature of the Milnor fiber of $X$. Then
$\lambda(Y)=\sigma(X)/8$
This relation, the famous “Casson invariant conjecture” has been extended in several directions [49].
@@
In a similar spirit, Curtis and Boden [4] proved an analogical relation for a version of $\mathop{SL}(2,\mathbb{C})$-Casson invariant $\lambda^{C}$ [16], with one-fourth of the full rank of the cohomology of the Milnor fiber (Milnor number).
Here we want to add two new observations to this story. Firstly, the $\mathop{SL}(2,\mathbb{C})$ relation extends to Seifert homology spheres of any number of fibers, using a new definition of Casson invariant $\lambda^{P}_{SL(2,\mathbb{C})}$ defined by Abouzaid and Manolescu [1] based on the work of Hamm [37] and Boden and Yokogawa [5]:
Proposition 1.5.
@@
Let $Y$ be a Seifert homology sphere and $X$ the corresponding complete intersection singularity with Milnor number $\mu(X)$. Then
$\lambda^{P}_{SL(2,\mathbb{C})}(Y)=\frac{\mu(X)}{4}.$
Secondly, for Seifert rational homology spheres, the equation (1.5) needs to be modified, bringing us back to one of our main topics - universal abelian covers. Namely, it is the spectrum of the universal abelian cover singularity, which needs to be considered to get $\lambda^{P}_{SL(2,\mathbb{C})}$. Moreover, to get a correct count, we need to quotient the spectrum by a certain action of $(\mathbb{Z}/2\mathbb{Z})^{2}$ (corresponding to the factor $1/4$ in (1.5)), which is not generally free in this case.
@@
If one asks what is the topological analog of the spectrum or the eigenvalues of the monodromy, the closest seem to be the “rotation numbers” of the components of the moduli space of connections. It is quite amusing that not only these numbers match (modulo the above action), but also the multiplicities in the spectrum match the Euler characteristics of the higher dimensional components of the moduli of connections in a way that when all combined together, the relation (1.5) holds. This suggests a deeper structure behind this numerical relation.
@@
The above relation between the monodromy of a singular surface with the moduli spaces of representations suggests a new nonabelian noncommutative Hodge Structure - a combination of Higgs bundles and Landau–Ginzburg theory. We sketch some examples and formulate a program for studying these structures. This more Hodge theoretic, rather than topological construction opens new ways of studying degenerations of nonabelian Hodge structures. We also propose a Landau-Ginzburg theory interpretation of $\widehat{Z}_{\sigma}(q)$. An additional novelty of this paper is that to the classical spectrum, we propose a derived spectrum. This spectrum has a separate Hodge theoretic merit. In addition, we conjecture that the derived spectrum produces invariants of 3-manifolds to which they are associated using Heegard splitting.

### v3

@@
[…]
$\widehat{Z}$ and Splice Diagrams
[…]
$\boldsymbol{\widehat{Z}}$ and Splice Diagrams
[…]
Sergei GUKOV a, Ludmil KATZARKOV b and Josef SVOBODA a
[…]
S. Gukov, L. Katzarkov and J. Svoboda
[…]
Received November 22, 2024, in final form August 16, 2025; Published online August 26, 2025
[…]
We study quantum $q$-series invariants of 3-manifolds $\widehat{Z}_{\sigma}$ of Gukov–Pei–Putrov–Vafa, using techniques from the theory of normal surface singularities such as splice diagrams. We show that the (suitably normalized) sum of all $\widehat{Z}_{\sigma}$ depends only on the splice diagram, and in particular, it agrees for manifolds with the same universal abelian cover. We use these ideas to find simple formulas for $\widehat{Z}_{\sigma}$ invariants of Seifert manifolds. Applications include a better understanding of the vanishing of the $q$-series $\widehat{Z}_{\sigma}$. Additionally, we study moduli spaces of flat $\operatorname{SL}_{2}(\mathbb{C})$ connections on Seifert manifolds and their relation to spectra of surface singularities, extending a result of Boden and Curtis for Brieskorn spheres to Seifert rational homology spheres with 3 singular fibers and to Seifert homology spheres with any number of fibers.
[…]
3-manifold topology; quantum invariant; surface singularity; splice diagram
[…]
57K31; 32S50; 32S25; 32S55
1 Introduction
@@
More specifically, we study $q$-series invariants $\widehat{Z}_{\sigma}(Y;q)$ of a 3-manifold $Y$, the so-called GPPV invariants, associated with quantum groups $\mathcal{U}_{q}(\mathfrak{g})$ at generic $|q|<1$ and labeled by a ${\rm spin}^{c}$ structure $\sigma\in{\rm spin}^{c}(Y)$ on $Y$. They are expected to have a categorification [34].
Mostly interested in topological and geometric aspects, throughout the paper, we consider the simplest non-trivial choice of $\mathfrak{g}=\mathfrak{sl}_{2}$, which corresponds to the “gauge group” $G=\operatorname{SU}(2)$, though much of the discussion can be generalized to higher rank root systems. [^1] We note that by “gauge group” one can mean at least two different things. One is the gauge group in complex Chern–Simons theory, for which $\widehat{Z}_{\sigma}(Y;q)$ provides a non-perturbative definition that behaves well under cutting and gluing. In contrast, the other one is a gauge group in 3d-3d correspondence and string theory fivebrane setup. The former group is a complexification of the latter, and it is the latter that we use to label $\widehat{Z}_{\sigma}(Y;q)$ in this paper. So, even when we talk about $G=\operatorname{SU}(2)$, the gauge group in complex Chern–Simons theory and its relations to quantum groups is, in fact, $\operatorname{SL}_{2}(\mathbb{C})$. Also, unless stated otherwise, throughout the paper, we work in the category of connected irreducible oriented negative definite plumbed 3-manifolds [50] with $b_{1}=0$, for which one can use the definition of $\widehat{Z}_{\sigma}(Y;q)$ as in [33]. (A reader interested in various generalizations and extensions is welcome to consult [2, 7, 29, 57, 58, 59, 61].)
Our main goal is to approach the invariants $\widehat{Z}_{\sigma}(Y;q)$ by using various methods of the singularity theory of normal complex surfaces, most notably the universal abelian covers and splice diagrams. Among other things, this new perspective leads to simpler expressions for $\widehat{Z}_{\sigma}(Y;q)$ for certain families of 3-manifolds (see Theorem 4.2) and sheds new light on what topological information they capture. Specifically, $\widehat{Z}_{\sigma}(Y;q)$ have close cousins $Z_{a}(Y;q)$ (without a hat!) that are not expected to admit a categorification, but are nevertheless more natural from the viewpoint of complex Chern–Simons theory [30]. While both sets of invariants can be defined independently, they are linearly related and, therefore, can be expected to contain roughly the same topological information about $Y$. By taking a closer look at the structure of the linear relation between $\widehat{Z}_{\sigma}(Y;q)$ and $Z_{a}(Y;q)$ through the looking glass of the singularity theory, we observe that $Z_{0}(Y;q)$ is, surprisingly, a much simpler invariant of $Y$ in the precise sense that we explain below.
@@
1.1 Main results
Let $Y$ be a closed, oriented, irreducible 3-manifold which is plumbed, see Section 2 for definitions. Moreover, assume that the plumbing matrix can be chosen to be negative definite. We call such a 3-manifold a negative definite plumbed manifold as is traditional. [^2] The reader should keep in mind that negative definiteness is a property of the plumbing graph rather than of $Y$. Note that these conditions are equivalent to $Y$ being a link of an isolated normal surface singularity $X$. Moreover, let us assume that $Y$ is a rational homology sphere, i.e., $b_{1}(Y)=0$, so it has finite $H:=H_{1}(Y,\mathbb{Z})$. Under these assumptions, $\widehat{Z}_{\sigma}(q)$ invariants were defined in [33].
The splice diagram is a combinatorial notion developed by Neumann, Eisenbud and Wahl [19, 54] for the study of singularities and 3-manifolds described above. Constructed from the plumbing graph, it contains certain essential part of it and proves very useful for studying of $\widehat{Z}_{\sigma}(q)$ invariants. The following theorem states that the series $Z_{0}(q)=\sum_{\sigma\in{\rm spin}^{c}(Y)}{\widehat{Z}_{\sigma}(q)}$ essentially depends on the splice diagram (see Section 3.1 for definitions). This applies in particular to homology spheres, where it provides a useful computational tool for obtaining the unique $\widehat{Z}_{\sigma}(q)=Z_{0}(q)$}. Denote by $\lambda(Y)$ the Casson–Walker invariant [69].
@@ 1.1 Main results
Topologically, splice diagrams determine universal abelian covers and vice versa [54, 60]. Universal abelian cover of a manifold $Y$ is the maximal cover of $Y$ with abelian deck group and by our assumptions, this cover is finite. We obtain the following corollary of the theorem.
@@
$q^{-6\lambda(Y_{1})}Z_{0}\bigl{(}Y_{1},q^{|H_{1}(Y_{1})|}\bigr{)}=q^{-6\lambda(Y_{2})}Z_{0}\bigl{(}Y_{2},q^{|H_{1}(Y_{2})|}\bigr{)}.$
For example, Seifert manifolds $Y_{1}=M(1;(7,1),(7,1),(7,4))$, $Y_{2}=M(1;(7,1),(7,2),(7,3))$ satisfy $Y_{1}^{\rm ab}=Y_{2}^{\rm ab}$ and $H=(\mathbb{Z}/7\mathbb{Z})^{2}$. Then $\lambda(Y_{1})=0$, $\lambda(Y_{2})=-21/2$ (see (3.4)) and
$Z_{0}\bigl{(}Y_{1},q^{49}\bigr{)}=q^{217/2}\bigl{(}1-3q^{35}+3q^{84}-q^{147}+q^{539}+\cdots\bigr{)}=q^{63}Z_{0}\bigl{(}Y_{2},q^{49}\bigr{)}.$
Seifert manifolds (Seifert fiber spaces) mentioned above are circle fibrations over 2-dimensional orbifolds. They admit “star-shaped” plumbing graphs and form basic building blocks of plumbed manifolds in Jaco–Shalen–Johannson decomposition. In this special case, we can go further with simplifications of $Z_{0}(Y)$ and we will give an explicit and rather simple formula, the “reduction” Theorem 4.2, using the data of Seifert fibration (rather than that of the plumbing graph). Using the group action of $H$ on $Y^{\rm ab}$, we will extract the separate $\widehat{Z}_{\sigma}(q)$ for each $\sigma$ in ${\rm spin}^{c}(Y)$.
The next theorem gives explicit formulas for $\widehat{Z}_{\sigma}(q)$ invariants of Seifert manifolds, see Section 4 for definitions and conventions. We introduce the “Laplace transform” $\mathcal{L}_{A}$, defined on monomials as $\mathcal{L}_{A}(t^{n})=q^{\frac{n^{2}}{4A}}$ and extended linearly to any formal power series in $t$. For a rational function $f(t)$, we denote $\mathrm{s.e.}\,f(t)$ the “symmetric expansion”, which is the average of the expansions of the given rational function as $t\rightarrow 0$ (as a Laurent power series in $t$) and as $t\rightarrow\infty$ (as a Laurent power series in $t^{-1}$). Finally, we pick a reference “anticanonical” ${\rm spin}^{c}{}$ structure $\sigma^{{\rm ac}}$ to express all $\widehat{Z}_{\sigma}(q)$ using the natural action of $H=H_{1}(Y,\mathbb{Z})=\langle g_{0},\dots,g_{k}\rangle$ on ${\rm spin}^{c}(Y)$ (see Sections 2.1 and 4 for notations and definitions).
@@
Let $Y=M(b;(a_{1},b_{1}),(a_{2},b_{2}),\dots,(a_{k},b_{k}))$ be a Seifert manifold over $S^{2}$ $(k\geq 3)$ with
$e=-b+\sum_{i=1}^{k}b_{i}/a_{i}<0.$
We set $A=a_{1}a_{2}\cdots a_{k}$ and $A_{i}=A/a_{i}$ for $i=1,\dots,k$. Then
$Z_{0}\bigl{(}q^{|H|}\bigr{)}=q^{\Lambda}\mathcal{L}_{A}(\mathrm{s.e.}\,f_{0}(t)),$ (1.2)
@@
$f_{0}(t)=\frac{\bigl{(}t^{A_{1}}-t^{-A_{1}}\bigr{)}\bigl{(}t^{A_{2}}-t^{-A_{2}}\bigr{)}\cdots\bigl{(}t^{A_{k}}-t^{-A_{k}}\bigr{)}}{\bigl{(}t^{A}-t^{-A}\bigr{)}^{k-2}},$
and $\Lambda=\Lambda(Y)$ is an explicit rational number given in (4.8). Moreover, for $h\in H$ we have
$\widehat{Z}_{h\sigma^{{\rm ac}}}(q)=q^{\Lambda/|H|}\mathcal{L}_{A|H|}(\mathrm{s.e.}\,f_{h\sigma^{{\rm ac}}}(t)),$
@@
$\sum_{h\in H}f_{h\sigma^{{\rm ac}}}(t)h=t^{(k-2)A-\sum_{i}A_{i}}\frac{\bigl{(}g_{1}t^{2A_{1}}-1\bigr{)}\cdots\bigl{(}g_{k}t^{2A_{k}}-1\bigr{)}}{\bigl{(}g_{0}t^{2A}-1\bigr{)}^{k-2}}.$
These formulae have both computational and conceptual significance. While in the original formulation of $\widehat{Z}_{\sigma}(q)$ we need to work with the plumbing data, which can be very large due to the presence of continued fractions, here we get $\widehat{Z}_{\sigma}(q)$ simply from the data of the Seifert fibration. It also sheds light on the role of the action of $H$, leading to a better understanding of the vanishing of the $q$-series for certain $\sigma$ in Corollary 4.3. Note that the independence of $f_{0}(t)$ on $b_{i}$ in (1.2) is an instance of Theorem 3.5.
1.2 Spectra in algebra and geometry
@@
The $\widehat{Z}_{\sigma}(Y)$ invariants are defined and studied using plumbing graphs. In singularity theory, negative definite plumbing graphs correspond to resolutions of the singularity which are smooth complex surfaces. Many deep results in singularity theory concern other smooth surfaces related to the singularity, namely those which are results of a smoothing. The theory of Milnor fibrations was originally developed for singularities of a single holomorphic function (hypersurface singularities) [43] and later extended to complete intersections [40].
@@
Spectrum of hypersurface singularity [67] was defined by Arnold, Steenbrink and Varchenko and generalized later by Steenbrink and Ebeling to complete intersections [18]. The spectrum is a collection of real numbers that refine the eigenvalues of the monodromy, using a natural mixed Hodge structure on the cohomology of Milnor fiber. Its most important property is the upper-semicontinuity under deformations.
@@
Recall that for pairwise coprime integers $p$, $q$, $r$, the Brieskorn homology sphere $Y=\Sigma(p,q,r)$ is the link of the Brieskorn singularity $X\colon x^{p}+y^{q}+z^{r}=0$. A relation between the Milnor fiber of $X$ and the topology of the link was given by Fintushel and Stern [24].
Theorem 1.4 ([24, Theorem 2.10]).
@@
Let $\lambda(Y)$ be the $\operatorname{SU}(2)$ Casson invariant of $Y$ and $\sigma(X)$ the signature of the Milnor fiber of $X$. Then $\lambda(Y)=\frac{\sigma(X)}{8}$.
@@
This relation, the celebrated ‘Casson invariant conjecture’ [53] has been extended in several directions, e.g., to the singularities of splice-type [49]. In a similar spirit, Curtis [16] defined the $\operatorname{SL}_{2}(\mathbb{C})$ Casson invariant $\lambda^{C}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)$ and with Boden proved a relation between $\lambda^{C}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)$ and the Milnor number $\mu(X)$, the rank of the cohomology of the Milnor fiber [4].
Theorem 1.5 ([4, Theorem 2.4.]).
With the notation above, $\lambda^{C}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)=\frac{\mu(X)}{4}$.
We extend this relation in two directions – to Seifert rational homology spheres with 3 singular fibers, and to Seifert homology spheres with any number of singular fibers.
For a Seifert rational homology sphere $Y=M(b;(a_{1},b_{1}),\dots,(a_{3},b_{3}))$, we need to consider the spectrum of the universal abelian cover singularity $X^{\rm ab}$, given by the equation $X^{\rm ab}\colon x_{1}^{a_{1}}+x_{2}^{a_{2}}+x_{3}^{a_{3}}=0$, where $a_{1}$, $a_{2}$, $a_{3}$ are not necessarily coprime. Instead of the Milnor number, we consider the number of orbits of a natural non-free $(\mathbb{Z}_{2})^{2}$-action of the spectrum $\operatorname{spec}\bigl{(}X^{\rm ab}\bigr{)}$ of $X^{\rm ab}$ (see (5.2) and (5.3)). The action replaces the factor of $1/4$ in Theorem 1.5.
On the topology side, we need the number $\lambda^{\rm nab}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)$ of characters of non-abelian flat connections, including the reducible ones (not present for Seifert homology spheres). This number was computed by Cui, Qiu and Wang [15].
Proposition 1.6 (Proposition 5.2).
Let $Y=M(b;(a_{1},b_{1}),\dots,(a_{3},b_{3}))$ be a Seifert rational homology sphere. With the notation above, we have
$\lambda^{\rm nab}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)=\left\lvert\frac{\operatorname{spec}\bigl{(}X^{\rm ab}\bigr{)}}{(\mathbb{Z}_{2})^{2}}\right\rvert$.
For Seifert homology spheres, we use an alternative definition of Casson invariant $\lambda^{P}_{\operatorname{SL}_{2}(\mathbb{C})}$ by Abouzaid and Manolescu [1], which takes into account higher-dimensional components of the moduli space of $\operatorname{SL}_{2}(\mathbb{C})$ flat connections on $Y$. Brieskorn singularity is replaced by the Brieskorn–Hamm complete intersection (see (5.7)).
Theorem 1.7 (Theorem 5.3).
@@
Let $Y$ be a Seifert homology sphere and $X$ the corresponding Brieskorn-Hamm complete intersection singularity with Milnor number $\mu(X)$. Then
$\lambda^{P}_{\operatorname{SL}_{2}(\mathbb{C})}(Y)=\frac{\mu(X)}{4}.$ (1.3)
The proof is merely a combination of formulas by Hamm [36] for Milnor number and Boden and Yokogawa [5] for the flat connections, but we could not find it in the literature. [^3] Boden and Curtis [4, p. 10] related the Euler characteristic of the union of the top-dimensional components of the moduli space with the Milnor number of a related hypersurface singularity., [^4] Recently, Muñoz derived the same result using gauge theory [44, Corollary B].
@@
If one asks what is the topological analogue of the spectrum or the eigenvalues of the monodromy, the closest seem to be the “rotation numbers” of the components of the moduli space of connections. In the course of the proof of Theorem 5.3, it is quite amusing to notice that not only these numbers match, but also the multiplicities of the spectral numbers match the Euler characteristic of the higher-dimensional components of the moduli space of connections. This suggests a deeper structure behind the numerical relation (1.3).
@@
The above relation between the monodromy of a singular surface with the moduli spaces of representations suggests the existence of new nonabelian noncommutative Hodge structure — a combination of Higgs bundles and Landau–Ginzburg theory, and related to it a new, derived spectrum. We sketch some examples and pose some questions about these structures. This more Hodge theoretic, rather than topological construction opens new ways of studying degenerations of nonabelian Hodge structures and a potential Landau–Ginzburg theory interpretation of $\widehat{Z}_{\sigma}(q)$.